# Learning How to Vote With Principles Axiomatic Insights Into the Collective Decisions of Neural Networks

### v1

@@ Discussion.
Moreover, for the WEC just three optimization objectives were enough to obtain the above competitive results. Since the MLP and CNN are not anonymous by design, they needed to optimize for anonymity as well. The MLP and CNN also needed to optimize for independence, while the WEC interestingly had enough implicit inductive bias toward satisfying independence—again highlighting non-trivial interference of the axioms and the network architecture.
@@
7 Discussion
With our axiomatic deep voting framework, we investigated the space of all voting rules by fruitfully combing voting theory and machine learning. The neural network explores the space and the voting-theoretic axioms evaluate the network, thus guiding the exploration. The universal approximation theorems (Hornik et al.,, 1989; Cybenko,, 1989) ensure that the neural networks are dense in the space of all voting rules, so all areas of that space can be explored with axiomatic deep voting. Arrow’s Impossibility Theorem (Arrow,, 1951) establishes insurmountable divisions of that space: e.g., the area of rules satisfying anonymity and Pareto does not intersect the area of rules satisfying independence.
@@ 7 Discussion
We tested a wide range of standard neural network architectures. However, future work could also investigate further architectures like Set Transformers, Graph Isomorphism Networks, or Deep Sets (which were used by Anil and Bao, (2021)) and, more generally, the transformer architecture (as a refinement our word embedding architecture). We also covered the most important voting-theoretic axioms, but yet more can be considered, e.g., monotonicitiy and transitivity (the latter then requires architectures that are not transitive by design). Finally, the large number of permutations causes a high statistical variance in testing the satisfaction of the independence axiom.
@@
Third, it seems promising to bridge notions of explainability in voting theory (Cailloux and Endriss,, 2016; Nardi et al.,, 2022; Boixel et al.,, 2022) and notions of explainability in AI (Adadi and Berrada,, 2018). In particular, is it possible to extract out a symbolic representation (e.g., in logic programming) of the rule that the model learned?
Fourth, studying the voting-theoretic concept of manipulatability via neural networks (Holliday et al.,, 2024) can be further connected to machine learning notions like adversarial attacks (Goodfellow et al.,, 2015) or performativity (Perdomo et al.,, 2020).
Fifth, from the point of view of geometric deep learning (Bronstein et al.,, 2021), axioms represent symmetries that the neural networks should learn. For example, anonymity says that the neural network should be invariant under the group action of the voter-permutation group on profiles; and neutrality says that the neural network should be equivariant under the group action of the alternative-permutation group on profiles and winning sets, respectively. It seems worth exploring this connection to geometric deep learning.
8 Conclusion
@@ Acknowledgments
For very helpful comments and discussions, we would like to thank Ben Armstrong, Balder ten Cate, Timo Freiesleben, Ronald de Haan, Alina Leidinger, and Christian List.
Appendix A List of all loss functions
We continue from section 4.4 and define the other loss functions that we use and discuss their differentiability. Recall that $\mathrm{KL}$ refers to Kullback–Leibler divergence. [^13] Though, in principle, other distance/similarity functions can be considered.
Anonymity
For convenience, we repeat the definition for the anonymity loss. Given the network $f_{w}$ and profile $\boldsymbol{P}$, uniformly sample $N$-many permutations $\pi_{1},\ldots,\pi_{N}$ of the set of voters of $\boldsymbol{P}$ and define
$L_{A}(f_{w},\boldsymbol{P}):=\frac{1}{N}\sum_{r=1}^{N}\mathrm{KL}\Big{(}f_{w}\big{(}e(\boldsymbol{P})\big{)},f_{w}\big{(}e(\pi_{r}(\boldsymbol{P}))\big{)}\Big{)}.$
Condorcet
If $\boldsymbol{P}$ has no Condorcet winner, $L_{C}(f_{w},\boldsymbol{P}):=0$, and otherwise, if that Condorcet winner is alternative $a$, define (recall $\overline{a}$ is the one-hot vector for alternative $a$)
Pareto
We define (recall that $\sigma$ is the sigmoid function and $n_{a\succ b}^{\boldsymbol{P}}=n$ means that all voters in $\boldsymbol{P}$ rank $a$ above $b$)
$L_{P}(f_{w},\boldsymbol{P}):=\sum_{a,b\text{ with }n_{a\succ b}^{\boldsymbol{P}}=n}\sigma\Big{(}f_{w}(e(\boldsymbol{P}))_{b}\Big{)}.$
Independence
Define $L_{I}(f_{w},\boldsymbol{P}):=0$ if $\boldsymbol{P}$ does not have at least two alternatives. Otherwise, randomly sample $N$-many pairs $(a_{r},b_{r})$ of distinct alternatives in $\boldsymbol{P}$ and randomly sample, for each ranking $P_{k}$ of $\boldsymbol{P}=(P_{1},\ldots,P_{n})$, a shuffling $P^{\prime}_{k}$ of $P_{k}$ in which, however, the order of $a_{r}$ and $b_{r}$ is the same as in $P_{k}$, and set $\boldsymbol{P}_{r}:=(P^{\prime}_{1},\ldots,P^{\prime}_{n})$. Write $\hat{y}:=f_{w}(e(\boldsymbol{P}))$ and $\hat{y}^{r}:=f_{w}(e(\boldsymbol{P}_{r}))$, and define
No winner
Writing $\hat{y}=f_{w}(e(\boldsymbol{P}))$ we want that at least one of the numbers in $p:=\big{(}\sigma(\hat{y}_{1}),\ldots,\sigma(\hat{y}_{m})\big{)}$ is above $0.5$, i.e., the maximum norm $\|p\|_{\infty}$ should be above $0.5$. Hence the more it is below that, the worse the loss:
$L_{NW}(f_{w},\boldsymbol{P}):=\max\big{(}0.5-\|p\|_{\infty},0\big{)}.$
For almost-everywhere differentiability use the distributivity of the differential operator over sums, the chain rule, and the almost-everywhere differentiability of the involved functions ($\mathrm{KL}$, $\sigma$, $\max$, $\|\cdot\|_{\infty}$).
@@ Appendix B Experiment 1
We add figure 7.
Plot for MLP with IC.
Plot for CNN with IC.
Plot for WEC with IC.
Plot for MLP with URN.
Plot for CNN with URN.
Plot for WEC with URN.
Plot for MLP with Mallows.
Plot for CNN with Mallows.
Plot for WEC with Mallows.
Plot for MLP with Euclidean.
Plot for CNN with Euclidean.
Plot for WEC with Euclidean.
Figure 7: Training the architectures (MLP, CNN, WEC; the three columns) on data from different voting rules (Plurality, Borda, Copeland; the three bars in each plot) using different sampling methods (IC, URN, MALLOWS, Euclidean; the four rows) and comparing the errors in both accuracy and axiom satisfaction.
@@ Appendix C Experiment 2
We add figure 8.
Plot for MLP with IC.
Plot for CNN with IC.
Plot for WEC with IC.
Plot for MLP with Mallows.
Plot for CNN with Mallows.
Plot for WEC with Mallows.
Figure 8: The top row is experiment 2 with IC sampling (repeated for convenience from the main text) and the bottom row is with Mallows sampling. For each architecture (columns from left to right: MLP, CNN, WEC), the accuracy and neutrality error are plotted across different ratios of augmented data. For example, $10\%$ in the $x$ axis means that the dataset consists of $10\%$ sampled data and $90\%$ augmented data.
@@ Appendix D Experiment 3
We add tables 3, 4, 5, 6, and 7.
@@
Table 3: The result of keeping on training a WEC model: each round adds 20k gradient steps to the previous one. Round 1 is with a learning rate of $10^{-3}$, round 2 with $10^{-4}$, round 3 with $5*10^{-5}$, and round 4 the same but with added optimization of independence. IC sampling.
Anon. Neut. Condorcet Pareto Indep. Average
Blacks 100 100 100 100 36.06 87.24
Stable Voting 100 100 100 100 39.74 87.96
Borda 100 100 94.12 100 35.32 85.9
Weak Nanson 100 100 100 100 39.62 87.92
Copeland 100 100 100 100 27.64 85.54
WEC n (NW, C, P) 100 100 95.4 100 43.88 87.88
Table 4: The average result over 5 runs with different seeds of experiment 3 for the ‘WEC n’ model and its closest rules. IC sampling.
Anon. Neut. Condorcet Pareto Indep. Average
Blacks 100 100 100 100 35.94 87.18
Stable Voting 100 100 100 100 40.8 88.14
Borda 100 100 94.36 100 35.36 85.94
Weak Nanson 100 100 100 100 40.26 88.06
Copeland 100 100 100 100 27.96 85.6
WEC n (NW, C, P) 100 100 94.66 100 41.94 87.34
Table 5: The average result over 5 runs with different seeds of experiment 3 for the ‘WEC n’ model and its closest rules. Mallows sampling.
@@
Table 6: Axiom satisfaction of different rules (top part of the table) and models (bottom part of the table). For Mallows. Rounded to one decimal. The names of the models are explained as follows: The letters after the architecture type indicated how the voting rule is computed from the model: p–plain (i.e., no averaging), n–neutrality-averaged, na-neutrality-and-anonymity-averaged. The letters in the brackets indicate which axioms the model optimized for during training: NW–No winner, A–Anonymity, C–Condorcet, P–Pareto, I–Independence. All models have been trained for 15k gradient steps with batch size 200 and the same seed. The WEC by far was the fastest.
WEC n
Stable Voting
Blacks
Borda
Weak Nanson
Copeland
WEC n 100 90.1 90.1 88.4 88 87.7
Stable Voting 100 96.22 91.57 93.28 92.23
Blacks 100 95.35 91.71 90.83
Borda 100 87.06 86.18
Weak Nanson 100 92.3

### v2

@@ Discussion.
Moreover, for the WEC just three optimization objectives were enough to obtain the above competitive results. Since the MLP and CNN are not anonymous by design, they needed to optimize for anonymity and independence as well. The WEC interestingly had enough implicit inductive bias toward satisfying independence—again highlighting non-trivial interference of the axioms and the network architecture.
@@
7. Discussion
With our axiomatic deep voting framework, we investigated the space of all voting rules by fruitfully combing voting theory and machine learning. The neural network explores the space and the voting-theoretic axioms evaluate the network, thus guiding the exploration. The universal approximation theorems (Hornik et al., 1989; Cybenko, 1989) ensure that the neural networks are dense in the space of all voting rules, so all areas of that space can be explored with axiomatic deep voting. Arrow’s Impossibility Theorem (Arrow, 1951) establishes insurmountable divisions of that space: e.g., the area of rules satisfying anonymity and Pareto does not intersect the area of rules satisfying independence.
@@ 7. Discussion
We tested a wide range of standard neural network architectures. However, future work could also investigate further architectures like Set Transformers, Graph Isomorphism Networks, or Deep Sets (which were used by Anil and Bao (2021)) and, more generally, the transformer architecture (as a refinement of our word embedding architecture). We also covered the most important voting-theoretic axioms, but yet more can be considered, e.g., monotonicitiy and transitivity (the latter then requires architectures that are not transitive by design). Finally, the large number of permutations causes a high statistical variance in testing the satisfaction of the independence axiom.
@@
Third, it seems promising to bridge notions of explainability in voting theory (Cailloux and Endriss, 2016; Nardi et al., 2022; Boixel et al., 2022) and notions of explainability in AI (Adadi and Berrada, 2018). In particular, is it possible to extract out a symbolic representation (e.g., in logic programming) of the rule that the model learned?
Fourth, studying the voting-theoretic concept of manipulatability via neural networks (Holliday et al., 2025) can be further connected to machine learning notions like adversarial attacks (Goodfellow et al., 2015) or performativity (Perdomo et al., 2020).
Fifth, from the point of view of geometric deep learning, [^22] See the work of Bronstein et al., 2021. Geometric Deep Learning: Grids, Groups, Graphs, Geodesics, and Gauges. arXiv:2104.13478. axioms represent symmetries that the neural networks should learn. For example, anonymity says that the neural network should be invariant under the group action of the voter-permutation group on profiles; and neutrality says that the neural network should be equivariant under the group action of the alternative-permutation group on profiles and winning sets, respectively. It seems worth exploring this connection to geometric deep learning.
8. Conclusion
@@ Acknowledgements.
For very helpful comments and discussions, we would like to thank Ben Armstrong, Balder ten Cate, Timo Freiesleben, Ronald de Haan, Thomas Icard, Alina Leidinger, Christian List, Ignacio Ojea, as well as the audiences at the ‘Workshop on Learning and Logic 2025’ at the University of Amsterdam and the ‘Social Choice for AI Ethics and Safety 2025’ workshop at AAMAS in Detroit. We are especially grateful to the JAIR editor Alessandro Farinelli and two anonymous reviewers.
Appendix A Hyperparameter Tuning
In this section, we investigate different hyperparameter choices for our models to motivate the choices we make in the main text.
A.1. Model Sizes
First, regarding the MLPs, Figure 7 tests the performance of different sizes. As mentioned in Section 5.4, we use the same sizes for our MLPs as Anil and Bao (2021): four hidden layers with 128 neurons each. To test this choice, we compare it to a smaller MLP with only two hidden layers with 128 neurons each, and to a larger MLP with six hidden layers with 128 neurons each plus layer norm. Figure 7 shows that all three MLPs achieve very similar accuracy. The larger MLP learns a bit more quickly than the other two, but it also has a higher variance in achieved accuracy. Hence the larger MLP does not yield a performance improvement. The results suggest that a smaller MLP might work, too, but, for continuity with the literature on the topic, when then choose their model sizes.
Second, regarding the CNNs and the WECs, the choice for the MLP size also dictates their sizes: in order to have a comparable capacity, they should have a roughly similar number of parameters. Indeed, with our choices of numbers and kinds of layers for the CNN and the WEC, we get, as described in Section 5.4, models that are roughly comparable in size.
[…]
Two plots describing the learning performance of different sized MLPs.
Figure 7. Different sized MLPs and their learning performance. The ‘small’ MLP has two hidden layers with 128 neurons each, the ‘standard’ MLP has four hidden layers with 128 neurons each (our and the literature’s choice), and the ‘large’ MLP has six hidden layers with 128 neurons each plus layer norm.
A.2. CNN Kernels and Pre-processing
Figure 8 investigates different choices for the hyperparameters of the CNN architecture. In Section 5.4, we described our choice: the two convolution layers have kernel sizes $(5,1)$ and $(1,5)$, respectively. Thus, the first kernel can pick up local patterns in the rankings of the voters, while the second kernel can pick up local patterns among the $i$-th preferred alternatives of the voters.
In image processing, ‘quadratic’ kernel sizes—e.g., (3,3)(3,3)—are more common, to pick up correlations of pixels with their surrounding pixels. In the voting setting, at least conceptually speaking, a quadratic surrounding does not make too much sense: Why should there be important ‘diagonal’ correlations, say between the $i$-th preferred alternative of voter $k$ and the $i+2$-th preferred alternative of voter $k+3$, especially if the voters should be permutable? On the contrary, vertical and horizontal correlations are important: Vertically, the first kernel captures patterns of correlation between a given alternative in a voter’s ranking and more or less preferred alternatives in that ranking; horizontally, the second kernel captures patterns of correlation between the $i$-th preferred alternative of a voter and the $i$-th preferred alternative of other voters.
Figure 8 shows that our choice of kernel size (top left) indeed achieves overall better results than a quadratic choice of kernel size (top right). Only for Condorcet and Independence, the quadratic choice is slightly better for some rules.
One might wonder, if one could still leverage diagonal correlations by first reordering the rankings of the voters in a profile so that ‘similar’ rankings are next to each other, before feeding the profile into the CNN. With such a pre-processing of the input, a quadratic kernel could be used since diagonal comparisons now make sense in such a similarity reordered profile. A standard way to formalize this notion of similarity is via Kendall Tau distance (as defined in footnote 8). We consider two versions of reordering a given profile $\boldsymbol{P}=(P_{1},\ldots,P_{n})$:
  • Global: Compute, for $k=2,\ldots,n$, the Kendall Tau distance $d_{k}$ between $P_{1}$ and $P_{k}$. Then reorder the profile starting with $P_{1}$ followed by the other rankings with ascending $d_{k}$. (In case of a tie, pick the ranking with minimal index first.)
  • Local: The reordered profile $\boldsymbol{P}^{\prime}=(P_{1}^{\prime},\ldots,P_{n}^{\prime})$ is computed recursively. Start with $P_{1}^{\prime}:=P_{1}$. Given $P_{k}^{\prime}$, we determine $P_{k+1}^{\prime}$ as follows. Go through the rankings that have not been picked yet (i.e., $\{P_{1},\ldots,P_{n}\}\setminus\{P_{1}^{\prime},\ldots,P_{k}^{\prime}\}$) and compute their Kendall Tau distance to $P_{k}^{\prime}$. Then $P_{k+1}^{\prime}$ is the ranking among these with the smallest Kendall Tau distance. (Again, tie-break via the indices.)
Figure 8 shows that adding either the global or the local version of Kendall Tau pre-processing overall does not reliably help the performance compared to our chosen setting (neither for our choice of kernel size nor for the quadratic choice). In some cases we do observe an improvement, as for example in the independence axiom satisfaction when learning the Copeland rule—however, this always comes with an additional loss, either in the accuracy or in the satisfaction of other axioms such as anonymity and neutrality.
[$(5,1)$, $(1,5)$, $32$, no KT] Refer to caption
[$(3,3)$, $(3,3)$, $32$, no KT] Refer to caption
[$(5,1)$, $(1,5)$, $32$, KT global] Refer to caption
[$(3,3)$, $(3,3)$, $32$, KT global] Refer to caption
[$(5,1)$, $(1,5)$, $32$, KT local] Refer to caption
[$(3,3)$, $(3,3)$, $32$, KT local] Refer to caption
[…]
Six barplots describing the performance of CNNs with different choices of hyperparameters.
Figure 8. Different hyperparameters of CNNs and their performance. The description [$(5,1)$, $(1,5)$, $32$, no KT] means that, for this CNN, the first kernel has size $(5,1)$, the second kernel has size $(1,5)$, the number of channels is 32, and the input is not Kendall Tau preprocessed. Similarly for the other descriptions. (The top left plot is our standard CNN setting and repeated from Figure 9 for convenience.)
A.3. Cross Validation
Finally, we corroborate our choice of hyperparameters by establishing their robust learning capabilities via cross validation in Table 3. For this, we IC-sample a fixed dataset of $100,000$ data points. We split the dataset into $10$ folds (each of size $10,000$). Looping over $k=0,\ldots,9$, we take fold $k$ as the test set and train the model on the data in the other $9$ folds for $8$ epochs. We record the achieved accuracy and loss (both on the training and the test set). Table 3 shows that, for all architectures with their chosen hyperparameters, we always get a high accuracy with little variance. This corroborates the robust learning capabilities of our architectures.
MLP
Testing fold number Train loss Train accuracy (in %) Test loss Test accuracy (in %)
0 0.013 97.7 0.031 95.1
1 0.013 97.7 0.033 94.8
2 0.012 97.9 0.032 95.4
3 0.009 98.5 0.031 95.4
4 0.018 96.9 0.038 94.4
5 0.018 96.9 0.04 94.1
6 0.025 95.8 0.042 93.7
7 0.011 97.9 0.03 95.1
8 0.013 97.7 0.029 95.9
9 0.008 98.7 0.025 96
Avg. 0.014 97.6 0.033 95
Std. dev. 0.005 0.8 0.005 0.7
CNN
Testing fold number Train loss Train accuracy (in %) Test loss Test accuracy (in %)
0 0.021 96.5 0.023 96.1
1 0.015 97.2 0.019 96.8
2 0.009 98.7 0.011 98.3
3 0.012 98 0.015 97.3
4 0.016 97.3 0.019 96.9
5 0.01 98.5 0.011 98
6 0.009 98.4 0.012 98
7 0.02 96.3 0.02 96.4
8 0.014 97.3 0.019 96.6
9 0.024 95.8 0.029 95.4
Avg. 0.015 97.4 0.018 97
Std. dev. 0.005 0.9 0.005 0.9
WEC
Testing fold number Train loss Train accuracy (in %) Test loss Test accuracy (in %)
0 0.005 99.4 0.006 99.5
1 0.005 99.7 0.005 99.8
2 0.006 99.6 0.006 99.5
3 0.007 99.3 0.008 99.1
4 0.035 96.3 0.037 95.8
5 0.004 99.8 0.004 99.9
6 0.012 98.6 0.014 98.5
7 0.01 99 0.011 98.7
8 0.01 98.5 0.009 98.6
9 0.011 98.4 0.011 98.3
Avg. 0.01 98.8 0.011 98.8
Std. dev. 0.009 1 0.009 1.1
Table 3. Cross validation of the three architectures on a dataset with $100,000$ data points IC-sampled from the Plurality rule with up to 55 voters and 5 alternatives. Training is for 8 epochs with a batch size of 200 (hence $8\times\frac{90,000}{200}=3,600$ gradient steps).
@@ Appendix B Experiment 1
We run the ‘correct for the right reasons’ experiment from Section 6.1 in more settings, reported in Figure 9 (part 1) and Figure 10 (part 2).
[…]
Six barplots describing further results of experiment 1 (part 1).
Figure 9. Part 1 of more settings of experiment 1 (Section 6.1). Varying architectures, rules, and sampling, while comparing the errors in both accuracy and axiom satisfaction.
[…]
Six barplots describing further results of experiment 1 (part 2).
Figure 10. Part 2 of more settings of experiment 1 (Section 6.1). Varying architectures, rules, and sampling, while comparing the errors in both accuracy and axiom satisfaction.
@@ Appendix C Experiment 2
We add further results to the ‘learning principles by example’ experiment from Section 6.2. Figure 11 and 12 show different choices of architecture, rules, and distribution for the first version of the experiment. Figure 13 and 14 show different choices of architecture, rules, and distribution for the second version of the experiment.
[…]
Six plots describing further results of experiment 2, in its first version.
Figure 11. The first version of experiment 2 (Section 6.2) with further architectures, rules, and distributions.
[…]
Two additional plots describing further results of experiment 2, in its first version.
Figure 12. The first version of experiment 2 (Section 6.2), but for the anonymity axiom (instead of neutrality). Since the WEC is anonymous by design, this can only be tested for MLP and CNN.
[…]
Six plots describing further results of experiment 2, in its second version.
Figure 13. The second version of experiment 2 (Section 6.2) with different distributions.
[…]
Three additional plots describing further results of experiment 2, in its second version.
Figure 14. The second version of experiment 2 (Section 6.2), but for the anonymity axiom (instead of neutrality). Since the WEC is anonymous by design, this can only be tested for MLP and CNN.
@@ Appendix D Experiment 3
We add further results on the ‘rule synthesis guided by principles’ experiment (Section 6.3).
D.1. More on the Experiment with IC
Regarding difference-making profiles, Figure 6 in the main text shows a profile where the model strongly disagrees with its 5 closest voting rules, i.e., the model’s winning set does not intersect the winning set of these rules. Figure 15 here shows a profile where the model weakly disagrees with all considered voting rules, i.e., the model’s winning set is not identical with the winning set of these rules. (For the main text profile, the model also happens to weakly disagree with all considered rules.) We found these profiles by going through 10,000 IC-sampled profiles and picking the weakly or strongly disagreeing profile with the smallest number of voters.
1 2 3 4 5 6 7
$a$ $b$ $c$ $e$ $b$ $b$ $d$
$e$ $d$ $e$ $c$ $c$ $c$ $e$
$d$ $c$ $a$ $d$ $a$ $a$ $b$
$b$ $a$ $d$ $a$ $d$ $d$ $c$
$c$ $e$ $b$ $b$ $e$ $e$ $a$
$\{b,c\}$ neutrality-averaged WEC, with sigmoids (rounded) aa:.38, bb:.54, cc:.50, dd:.39, ee:.39
$\{b\}$ Plurality, Weak Nanson, Kemeny-Young
$\{c\}$ Borda, Copeland, Llull, Blacks, Coombs
$\{d\}$ Anti-Plurality, Baldwin
$\{e\}$ Instant Runoff TB
$\{b,d\}$ Stable Voting
$\{b,c,d\}$ Uncovered Set, Banks
$\{b,c,e\}$ PluralityWRunoff PUT
$\{a,b,c,d,e\}$ Top Cycle
[…]
A profile on which the ‘WEC n’ model weakly disagrees (IC sampling).
Figure 15. IC sampling: Profile where the ‘WEC n’ model weakly disagrees (i.e. non-identical winning sets) with existing voting rules.
Table 4 suggests that further optimization does not further improve axiom satisfaction.
@@ D.1. More on the Experiment with IC
Table 4. IC sampling: The result of keeping on training an WEC model. Each round adds 20k gradient steps to the previous one. Round 1 is with a learning rate of $10^{-3}$, round 2 with $10^{-4}$, round 3 with $5*10^{-5}$, and round 4 the same but with added optimization of independence.
Table 5 shows the statistical robustness of the axiom satisfaction achieved by the model.
Anon. Neut. Condorcet Pareto Indep. Avg.
Blacks 100 100 100 100 36.04 87.2
Stable Voting 100 100 100 100 40.48 88.1
Borda 100 100 93.82 100 37.72 86.32
Weak Nanson 100 100 100 100 38.28 87.68
Copeland 100 100 100 100 28.54 85.72
WEC n (NW, C, P) 100 100 96.78 100 45.9 88.54
Table 5. IC sampling: Take the average over 5 runs of checking the axiom satisfaction of the ‘WEC n’ model and its closest rules.
Figure 16 shows the evolution of the semantic losses during optimization and the run times, showing that the WEC not only achieves the best results but also does so most quickly.
[…]
Three plots describing the loss evolution during experiment 3.
Figure 16. The loss evolution during experiment 3. The run times of optimization were as follows: 5h 29min (MLP), 6h 08min (CNN), 2h 05min (WEC).
D.2. The Experiment with Mallows
Table 6 shows the result of the experiment from Section 6.3 but with Mallows sampling (instead of IC sampling), using the parameters discussed in Section 5.1. Table 7 shows how similar the best model is to its closest rules. Figure 17 presents a profile where the model differs from all considered rules. Table 8 shows the statistical robustness of the axiom satisfaction achieved by the model.
@@ D.2. The Experiment with Mallows
Table 6. Mallows sampling: Axiom satisfaction of different rules (top part of the table) and models (bottom part of the table). Otherwise like Table 1 from the main text.
Identity accuracy WEC n Stable Voting Blacks Borda Weak Nanson Copeland
WEC n 100 89.1 89.4 88.3 87.3 87.2
Stable Voting 100 95.61 91.04 93.47 92.16
Blacks 100 95.43 91.71 90.82
Borda 100 87.14 86.25
Weak Nanson 100 92.08